200 graduating students are selected and it is found that 114 will be awarded the first class honors degrees calculate the 95% confidence interval for the proportion of the graduating students who will receive first class honors degrees

1
Expert's answer
2022-06-06T23:56:59-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

"\hat{p}=\dfrac{X}{N}=\dfrac{114}{200}=0.57"

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},""\hat{p}+z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})""=(0.57-1.96\times \sqrt{\dfrac{0.57(1-0.57)}{200}},""0.57+1.96\times \sqrt{\dfrac{0.57(1-0.57)}{40}})""=(0.5014, 0.6386)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.5014 < p < 0.6386," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.5014, 0.6386)."